# Computing the Matching Distance of $2$-Parameter Persistence Modules from Critical Values

### v2

@@ Abstract
The exact computation of the matching distance for multi-parameter persistence modules is an active area of research in computational topology. Achieving an easily obtainable exact computation of this distance would allow multi-parameter persistent homology to be a viable option for data analysis. In this paper, we provide theoretical results for the computation of the matching distance in two dimensions along with a geometric interpretation of the lines through parameter space realizing this distance. In particular, we show that it is possible to avoid considering vertical and horizontal lines. To prove this, lines with slope 1 play an essential role.
@@ 1 Introduction
To compute the matching distance between two $n$-parameter persistence modules, one uses the fact that restricting an $n$-parameter filtration to any affine line of positive slope through the parameter space yields a $1$-parameter filtration. There is therefore a corresponding restriction of the $n$-parameter persistence module to a $1$-parameter module along that line. This construction allows for the knowledge and computational methods from the $1$-dimensional case to be applied to the $n$-dimensional case.
@@
where $\hat{m}^{L}$ is a normalization term explained in Section 2. However, exactly computing this distance is not an easy task given the nature of its definition. As a first step towards an exact computation, several approximations of this distance have been provided for $2$-dimensional persistence modules [4, 22]. The exact computation of the matching distance is currently only possible for $2$-dimensional modules [20], with recent computational improvements in terms of time complexity [5]. These methods exploit the duality of points and lines in the plane.
@@
While approximated computations can be very useful in practice, computing the exact value of the matching distance has its own merits. One of them is that it can distinguish modules for which the rank invariant is a complete invariant, such as rectangle decomposable modules, as the matching distance is a metric for the rank invariant. [7]
@@
In this paper, we propose a method to compute the matching distance based on a refinement of the framework developed in [1], in which we compute the rank invariant of a multi-persistence module from a finite set of critical parameter values. These critical parameter values capture all the changes in homology occurring throughout the multi-filtration. They may also be used to partition the set of positive lines in the parameter space into equivalence classes, where each equivalence class maintains the same birth and death ordering within the restricted module for all possible homology classes.
@@
Based on these results, we formulate the idea that the critical values must be relevant to the choice of lines for the computation of the matching distance. Leveraging the framework of [1] to compute the matching distance allows us to reduce the number of lines necessary to compute the matching distance to a finite set, thus reducing the computation to a maximum rather than a supremum without exploiting the point-line duality used in [20]. Through some examples, we show that considering only lines passing through pairs of critical values is not sufficient. This is because the definition of matching distance depends on the bottleneck distance, and lines in the same equivalence class might not be such that the bottleneck distance is always achieved by the same pairing of births/deaths, as will be further explained in Section 3.
@@
To overcome this problem, we analyze where switches might happen in the matching which achieves the bottleneck distance, identifying a set of points in the parameter space which we later on refer to as the set of switch points. This set allows us to refine our equivalence relation on the set of positive lines (see Section 4) by partitioning lines through the parameter space using the union of the critical parameter values with the switch points. Using this union, it is possible to identify all the lines at which the matching distance can be realised. We explain in detail how to compute all relevant parameter values and show that the matching distance is attained either at: a line through a pair of points in this union, or a line of diagonal slope through exactly one of the points in the union (see Theorem 3.1).
The contribution of this paper is to advance the state of the art, although still restricted to two dimensions, in two ways: computing the matching distance in a way which is both geometrically interpretable and implementable. In contrast to other methods such as [20, 5], we provide a geometric understanding of different lines, including horizontal, vertical, and diagonal lines, as well as lines through critical values, and their contribution to the matching distance. In addition, the method we propose leads to explicit computations of important lines, which will lead to an algorithm that can be implemented.
@@ 2.1 Notation and definitions
$I[b,d)_{s}=\left\{\begin{array}[]{cc}\mathbb{F}&\text{if }b\leq s<d\\ 0&\text{otherwise}\end{array}\right.,\quad I[b,d)(s\leq t)=\left\{\begin{array}[]{cc}\mathrm{id_{\mathbb{F}}}&\text{if }b\leq s\leq t<d\\ 0&\text{otherwise}\end{array}\right.$
@@
Let $M$, $N$ be two $1$-parameter persistence modules, with persistence diagrams $\operatorname{dgm}M$ and $\operatorname{dgm}N$.
@@
By definition, the matching distance is computed by taking the supremum of the bottleneck distance over all (infinitely many) lines with positive slope through the parameter space. However, in [1], we have shown that this set of lines may be partitioned into a finite number of equivalence classes. For the two-parameter case (i.e., for $n=2$), we will use these equivalence classes to generate a finite set of lines from which one may compute the matching distance. Therefore, we introduce notation and definitions necessary to understand these equivalence classes below.
Henceforth in this paper, we always assume $n=2$, which is the case for which we have extremely explicit results. We introduce the following definitions (with associated figures) following [1].
@@
$S_{A}(u):=\begin{cases}\{u\}&\textrm{ if }A=\{1,2\}\ ,\\ \{(x_{1},x_{2})\in\mathbb{R}^{2}\ |\ x_{1}=u_{1},x_{2}>u_{2}\}&\textrm{ if }A=\{1\}\ ,\\ \{(x_{1},x_{2})\in\mathbb{R}^{2}\ |\ x_{1}>u_{1},x_{2}=u_{2}\}&\textrm{ if }A=\{2\}\ .\\ \end{cases}$
@@
In this paper, it is necessary for us to extend this equivalence relation to include points in the projective completion of the real plane, $\mathbb{P}^{2}=\mathbb{R}^{2}\cup\ell_{\infty}$, with $\ell_{\infty}$ the line at infinity of $\mathbb{P}^{2}$. Points on the line at infinity are given by homogeneous coordinates $[x_{0}:x_{1}:x_{2}]$ with $x_{0}=0.$ Note that a line with positive slope in $\mathbb{R}^{2}$ will intersect $\ell_{\infty}$ at exactly one point $[0:v_{1}:v_{2}]$ with $\vec{v}=(v_{1},v_{2})\succ 0$ giving the direction of the line.
@@
$S_{A}(v):=\begin{cases}\{v\}&\textrm{ if }A=\{1,2\}\ ,\\ \{[0:x_{1}:x_{2}]\in\ell_{\infty}\ |\ x_{1}=v_{1},x_{2}>v_{2}\}&\textrm{ if }A=\{1\}\ ,\\ \{[0:x_{1}:x_{2}]\in\ell_{\infty}\ |\ x_{1}>v_{1},x_{2}=v_{2}\}&\textrm{ if }A=\{2\}\ .\\ \end{cases}$
@@
$L\sim_{U}L^{\prime}\text{ if and only if }A^{L}_{u}=A^{L^{\prime}}_{u}\;\forall u\in U\ .$
@@
$p_{L}(u)-p_{L}(v)=\left\{\begin{array}[]{ll}\operatorname{push}_{L}(u)_{2}-\operatorname{push}_{L}(v)_{2}&\mbox{ if $L$ has slope greater than 1,}\\ \operatorname{push}_{L}(u)_{1}-\operatorname{push}_{L}(v)_{1}&\mbox{ otherwise.}\end{array}\right.$ (5)
@@ 2.2 Working assumptions
Property (P).
There exists a finite set $C_{M}$ of points in $\mathbb{R}^{n}$, called critical values, for which the ranks of transition maps in the $n$-parameter persistence module $M$ are completely determined by the ranks of transition maps between values in $\overline{C}_{M}$, the closure of $C_{M}$ under least upper bound. More specifically, for any $u,v\in\mathbb{R}^{n}$,
@@
As we will see in Lemma 3.2 later, these bijections will extend to a bijection, denoted $\Gamma_{L,L^{\prime}}$, between the sets of matchings that are used to compute both $d_{B}(M^{L},N^{L})$ and $d_{B}(M^{L^{\prime}},N^{L^{\prime}})$ Moreover, the bijection $\Gamma_{L,L^{\prime}}$ can be used to compute the cost of a matching for line $L^{\prime}$ from the cost of the corresponding matching for line $L$, provided the lines are in the same equivalence class.
Ideally, one would like to use $\Gamma_{L,L^{\prime}}$ to show that a line $L$ on which $d_{B}(M^{L},N^{L})$ attains a maximum is a line which passes through two points in $\overline{C}_{M}\cup\overline{C}_{N}.$ However, this is not the case, as Examples 2.11 and 2.13 show. The reason is that, within an $\sim_{\overline{C}_{M}\cup\overline{C}_{N}}$-equivalence class, it is possible that there is no “winning” matched pair of points or “winning” matching that works for all lines within that equivalence class, i.e., the winning matching may “switch” within the same $\sim_{\overline{C}_{M}\cup\overline{C}_{N}}$ equivalence class.
@@
[…]
@@ 2.3 Examples
We now give some simple examples of computation of the matching distance in the case of two parameters. The considered persistence modules will be direct sums of rectangle modules, i.e. persistence module that are trivial outside a supporting rectangle $[u_{1},v_{1})\times[u_{2},v_{2})$, with $u=(u_{1},u_{2})$, $v=(v_{1},v_{2})\in\mathbb{R}^{2}$, and equal to $\mathbb{F}$ with all the transition maps equal the identity on the supporting rectangle with bottom left corner $u$ and a top right corner $v$.
@@
Let $M$ be the rectangle module with support $[2,\infty)\times[2,7)$, and let $N$ be the rectangle module with support $[2,\infty)\times[2,10)$.
Figure 3: The persistence module considered in Example 2.11: The matching distance is not attained at any line through two critical points.
For lines $L$ with equation $y=mx+q$ with $m\leq 1$ and $q\geq 2$, we get weighted bottleneck distance $5$, thus achieving the matching distance, because the weighted distance is smaller than $5$ for all other lines. In particular, the line through $(2,2)$, point in $C_{M}\cap C_{N}$, with diagonal direction achieves the matching distance between $M$ and $N$.
Symmetrically, if we consider the rectangle persistence modules $M=[2,7)\times[2,\infty)$ and $N=[2,10)\times[2,\infty)$, the situation is reversed, that is, the matching distance is realised by lines of slope greater or equal than $1$.
@@
Let $M$ be the rectangle decomposable module given as the direct sum of two rectangle modules, one with support $[0,7)\times[0,7)$, the other one with support $[0,7)\times[4,11)$. Let $N$ be the rectangle decomposable module given as the direct sum of the rectangle modules with support $[0,7)\times[0,11)$ and $[0,7)\times[4,7)$.
For diagonal lines $L$ with equation $y=x+q$ for $q\leq 0$ or $q\geq 4$, we have that $M^{L}=N^{L}$ so that $d_{B}(M^{L},N^{L})=0$. For $40<q<$ 4, a straightforward computation shows that the maximum bottleneck distance achieved is for $q=2$, where the cost is $2$. However, for the line $L^{\prime}$ through the points $(0,0)$ and $(7,11)$, the bottleneck distance is $d_{B}(M^{L},N^{L})=4$ and $\hat{m}^{L^{\prime}}=\frac{7}{11}$, so that the matching distance between $M$ and $N$ is at least $\frac{28}{11}$. This example shows that diagonal lines do not suffice.
@@
Let $M$ be the rectangle decomposable module which is the direct sum of two rectangle modules, one with underlying rectangle $[0,7)\times[0,8)$, the other one with underlying rectangle $[0,7)\times[4,11)$. Let $N$ be the rectangle decomposable module sum of two rectangle modules whose underlying rectangles are $[0,7)\times[4,8)$ and $[0,7)\times[0,11)$.
@@
Figure 6: Weighted bottleneck distance for the modules of Example 2.13 across the line parameters. On the horizontal axis we have the values of the angle $\theta$, giving the slope of the line. On the vertical axis we have the parameter $b$ giving the origin of the line. The red cross corresponds to the maximum value over all lines computed in the approximation, the black cross the maximum value for lines through pairs of critical points and the blue cross is the maximum value for diagonal lines. This particular example shows that it is not sufficient to consider only lines through pairs of critical points or diagonal lines. This is because the matching distance must be at least as much as the value computed at the red cross.
The following considerations confirms this: for diagonal lines $L$ with equation $y=x+q$ for $q\leq 1$ or $q\geq 4$, we have that $M^{L}=N^{L}$ so that $d_{B}(M^{L},N^{L})=0$. For $41<q<$ 4, straightforward computation shows that the maximum bottleneck distance achieved is for $q=2.5$, where the cost is $\frac{3}{2}$. However, for the line $L^{\prime}$ through the points $(\frac{7}{2},6)$ and $(7,11)$, the bottleneck distance is $d_{B}(M^{L},N^{L})=3$ and $\hat{m}^{L^{\prime}}=\frac{7}{10}$, so that the matching distance between $M$ and $N$ is at least $\frac{21}{10}$. Out of the lines through two points in the closure with respect to lowest upper bound of $C_{M}\cup C_{N}$, the only line that does not give $M^{L}=N^{L}$ is the line through $(0,0)$ and $(7,11)$, which has weighted bottleneck distance $\frac{21}{11}$. This example shows that it does not suffice to consider only diagonal lines and lines through two points in the closure with respect to the lowest upper bound of $C_{M}\cup C_{N}$.

### v3

@@ Abstract
The exact computation of the matching distance for multi-parameter persistence modules is an active area of research in computational topology. Achieving an easily obtainable exact computation of this distance would permit multi-parameter persistent homology to be a viable option for data analysis. For this purpose, two approaches are currently available, limited to persistence with parameters from $\mathbb{R}^{2}$: authors of [25, 7] work in the discrete setting and apply the point-line duality; authors of [18, 19] work in the smooth setting while remaining in the primal plane. In this paper, we streamline the computation of the matching distance in the combinatorial setting while staying in the primal plane. In doing so, besides connecting results from the literature, we give explicit formulas for the switch points needed by all the available methods and we show that it is possible to avoid considering vertical and horizontal lines. For the latter, lines with slope 1 play an essential role.
@@ 1 Introduction
To compute the matching distance between two $n$-parameter persistence modules, one uses the fact that restricting an $n$-parameter filtration to any affine line $L$ of positive slope through the parameter space yields a $1$-parameter filtration. There is therefore a corresponding restriction of the $n$-parameter persistence module $M$ to a $1$-parameter module $M^{L}$ along that line. This construction allows for the knowledge and computational methods from the $1$-dimensional case to be applied to the $n$-dimensional case.
@@
where $\hat{m}^{L}$ is a normalization term explained in Section 2. However, exactly computing this distance is not an easy task given the nature of its definition. As a first step towards an exact computation, several approximations of this distance have been provided for $2$-dimensional persistence modules [5, 16, 27].
@@
While approximated computations can be very useful in practice, computing the exact value of the matching distance has its own merits. One of them is that it can distinguish modules for which the rank invariant is a complete invariant, such as rectangle decomposable modules, as the matching distance is a metric for the rank invariant [9].
Methods for the exact computation of the matching distance are currently only available for $2$-dimensional modules. One approach, carried out in [18, 19], is dedicated to smooth input data and is based on the concept of an extended Pareto grid. Another approach, carried out in [25], is dedicated to combinatorial input and is based on looking at lines as points in the dual plane. The paper [7] presents improvements to this approach in terms of time complexity, still exploiting the point-line duality in the plane.
In this paper, we streamline the computation of the matching distance for combinatorial input data in the primal plane. Our primary goal for this is to understand the role of lines with special slopes like diagonal, vertical, and horizontal lines. As a secondary goal, we aim to prove that the matching distance in the combinatorial setting is realized as a maximum. In doing so, we also re-interpret results from different papers using the primal plane as a reference.
@@
The present work is based on a refinement of the framework developed in [1], in which we compute the rank invariant of a multi-persistence module from a finite set of critical parameter values. These critical parameter values capture all the changes in homology occurring throughout the multi-filtration. They may also be used to partition the set of positive lines in the parameter space into equivalence classes, where each equivalence class maintains the same birth and death ordering within the restricted module for all possible homology classes.
@@
Based on these results, one sees that the critical values must be relevant to the choice of lines for the computation of the matching distance. Leveraging the framework of [1] to compute the matching distance allows us to reduce the number of lines necessary to compute the matching distance to a finite set, thus reducing the computation to a maximum rather than a supremum without exploiting the point-line duality used in [25]. Through some examples, we show that considering only lines passing through pairs of critical values is not sufficient. This is because the definition of matching distance depends on the bottleneck distance, and lines in the same equivalence class might not be such that the bottleneck distance is always achieved by the same pairing of births/deaths, as will be further explained in Section 3.
@@
To overcome this problem, we analyze where switches might happen in the matching that achieves the bottleneck distance, identifying a set of points in the parameter space that we later on refer to as the set of switch points. This set allows us to refine our equivalence relation on the set of positive lines (see Section 4) by partitioning lines through the parameter space using the union of the critical parameter values with the switch points. Using this union, it is possible to identify all the lines at which the matching distance can be realised.
We explain in detail how to compute all relevant parameter values in the primal plane and show that the matching distance is attained either at: a line through a pair of points in this union, or a line of diagonal slope through exactly one of the points in the union (see Theorem 3.1). Where appropriate, we also detail the connections and differences between related papers.
In conclusion, the novelty of this paper is to provide: a geometric understanding of different lines, including horizontal, vertical, and diagonal lines, and their contribution to the matching distance; explicit formulas for the switch points; a common framework where existing approaches in the literature can be compared.
@@ 2.1 Notation and definitions
$I[b,d)_{s}=\left\{\begin{array}[]{cc}\mathbb{F}&\text{if }b\leq s<d\\ 0&\text{otherwise}\end{array}\right.,\quad I[b,d)(s\leq t)=\left\{\begin{array}[]{cc}\mathrm{id_{\mathbb{F}}}&\text{if }b\leq s\leq t<d\\ 0&\text{otherwise.}\end{array}\right$.
@@
Let $M$, NN be two $1$-parameter persistence modules, with persistence diagrams $\operatorname{dgm}M$ and $\operatorname{dgm}N$.
@@
By definition, the matching distance is computed by taking the supremum of the bottleneck distance over all (infinitely many) lines with positive slope through the parameter space. However, in [1], we have shown that this set of lines may be partitioned into a finite number of equivalence classes. For the two-parameter case (i.e., for $n=2$), we will use these equivalence classes to generate a finite set of lines from which one may compute the matching distance.
Henceforth in this paper, we always assume $n=2$, which is the case for which we have explicit results.
We first review the notation and definitions (with associated figures) necessary to understand these equivalence classes following [1] for the two-parameter case. We summarize the relationship with analogous notions in the literature in Section 2.2.
@@
In this paper, it is necessary for us to extend this equivalence relation to include points in the projective completion of the real plane, $\mathbb{P}^{2}=\mathbb{R}^{2}\cup\ell_{\infty}$, with $\ell_{\infty}$ the line at infinity of $\mathbb{P}^{2}$. Points on the line at infinity are given by homogeneous coordinates $x_{0}:x_{1}:x_{2}]$ with $x_{0}=0.$ Note that a line with positive slope in $\mathbb{R}^{2}$ will intersect $\ell_{\infty}$ at exactly one point $0:v_{1}:v_{2}]$ with $\vec{v}=(v_{1},v_{2})\succ 0$ giving the direction of the line.
@@
[…]
@@
LUL $\sim_{U}L^{\prime}\text{ if and only if }A^{L}_{u}=A^{L^{\prime}}_{u}\;\forall u\in$ U\ .
@@
[…]
For sufficiently nice modules (such as those that satisfy Property (P) from Section 2.3), there is a finite set $C$ of points in $\mathbb{R}^{n}$ with the property that, for a given line $L$, the set $\left\{\operatorname{push}_{L}(c)|c\in C\right\}$ can be used to determine the coordinates of the points in the persistence diagram associated to $L$. (See, for example, [1, Thm. 2])
Lemma 2.10.
Proof.
More explicitly, the multi-bijection $\gamma_{M}$ behaves as follows.
Similarly, there is some $j\in\{1,2\}$ such that
2.2 Connections to related works
As mentioned in Section 1, there are other papers that look to understand the matching distance for bi-persistence modules. Our approach to computing the matching distance in this paper is closely related to two groups of papers:
  • The papers [25, 31] present an approach in the dual space of lines, and
  • The papers [15, 16, 18] present an approach in topological persistence.
Where appropriate throughout the paper, we will highlight the similarities and differences between our paper and these two groups (termed the “dual approach” and the “topological approach”, respectively). The first such section here will provide connections and analogues to the dual and topological approach for the content in Section 2.1.
Firstly, note that Definition 2.8 is related to the construction of the extended Pareto grid of topological persistence from [15]. One of the main differences between this paper and [15] is the fact that we assume we have a finite number of critical parameter values which determine the persistence diagrams along lines, while the persistence diagram along a line $L$ in [15] is determined by the intersection of $L$ with the extended Pareto grid (see Theorem 2 in [15]).
Secondly, we note that the bijection $\Gamma_{L,L^{\prime}}$ from Lemma 2.10 corresponds to the transport of matchings introduced in [15] (Subsection 4.1). Specifically, the map $\Gamma_{L,L^{\prime}}$ associates the matching corresponding to the line $L$ with the matching corresponding to the line $L^{\prime}$, constructed by tracking the movement of points in the persistence diagrams as $L$ transitions to $L^{\prime}$. A similar concept also appears in [25], in the proof of Theorem 4.6. In all three papers ([25, 15] and this paper) it is necessary to identify a bijection between points in the persistence diagrams along $L$ and $L^{\prime}$, in order to bound the weighted cost of the bottleneck distance on some equivalence class of lines. This bound and the finiteness of the number of such equivalence classes is what allows us to obtain a maximum for the matching distance.
@@ 2.3 Working assumptions
Property.
[P] There exists a finite set $C_{M}$ of points in $\mathbb{R}^{n}$, called critical values, for which the ranks of transition maps in the $n$-parameter persistence module $M$ are completely determined by the ranks of transition maps between values in $\overline{C}_{M}$, the closure of $C_{M}$ under least upper bound. More specifically, for any $u,v\in\mathbb{R}^{n}$,
@@
As we will see in Lemma 2.10 later, these bijections will extend to a bijection, denoted $\Gamma_{L,L^{\prime}}$, between the sets of matchings that are used to compute both $d_{B}(M^{L},N^{L})$ and $d_{B}(M^{L^{\prime}},N^{L^{\prime}})$ . Moreover, the bijection $\Gamma_{L,L^{\prime}}$ can be used to compute the cost of a matching for line $L^{\prime}$ from the cost of the corresponding matching for line $L$, provided the lines are in the same equivalence class.
Ideally, one would like to use $\Gamma_{L,L^{\prime}}$ to show that a line $L$ on which $d_{B}(M^{L},N^{L})$ attains a maximum is a line which passes through two points in $\overline{C}_{M}\cup\overline{C}_{N}.$ However, this is not the case, as Examples 2.12 and 2.14 show. The reason is that, within an $\sim_{\overline{C}_{M}\cup\overline{C}_{N}}$-equivalence class, it is possible that there is no “winning” matched pair of points or “winning” matching that works for all lines within that equivalence class, i.e., the winning matching may “switch” within the same $\sim_{\overline{C}_{M}\cup\overline{C}_{N}}$ equivalence class.
@@
Finally, because Example 2.12 shows that the diagonal direction is also needed, we set $\overline{\Omega}$ to be the union of $\Omega$ with the singleton $\{[0:1:1]\}$. This will allow for the computation of the matching distance between $M$ and $N$ from lines through two points in $\overline{C}_{M}\cup\overline{C}_{N}\cup\overline{\Omega}$ via Theorem 3.1, assuming that $M$ and $N$ satisfy properties (P) and (Q).
@@ 2.4 Examples
We now give some simple examples of computation of the matching distance in the case of two parameters. The considered persistence modules will be direct sums of rectangle modules, i.e. persistence module that are trivial outside a supporting rectangle $u_{1},v_{1})\times[u_{2},v_{2})$, with $u=(u_{1},u_{2})$, $v=(v_{1},v_{2})\in\mathbb{R}^{2}$, and equal to $\mathbb{F}$ with all the transition maps equal the identity on the supporting rectangle with bottom left corner $u$ and a top right corner $v$.
@@
Let $M$ be the rectangle module with support $2,\infty)\times[2,7)$, and let $N$ be the rectangle module with support $2,\infty)\times[2,10)$.
Figure 3: The persistence module considered in Example 2.12: The matching distance is not attained at any line through two critical values.
For lines $L$ with equation $y=mx+q$ with $m\leq 1$ and $q\leq 7-2m$, we get weighted bottleneck distance $3$, thus achieving the matching distance, because the weighted distance is smaller than $3$ for all other lines. In particular, the line through $(2,2)$, point in $C_{M}\cap C_{N}$, with diagonal direction achieves the matching distance between $M$ and $N$.
Symmetrically, if we consider the rectangle persistence modules $M=[2,7)\times[2,\infty)$ and $N=[2,10)\times[2,\infty)$, the situation is reversed, that is, the matching distance is realised by lines of slope greater than or equal to $1$.
@@
Let $M$ be the rectangle decomposable module given as the direct sum of two rectangle modules, one with support $0,7)\times[0,7)$, the other one with support $0,7)\times[4,11)$. Let $N$ be the rectangle decomposable module given as the direct sum of the rectangle modules with support $0,7)\times[0,11)$ and $0,7)\times[4,7)$.
For diagonal lines $L$ with equation $y=x+q$ for $q\leq 0$ or $q\geq 4$, we have that $M^{L}=N^{L}$ so that $d_{B}(M^{L},N^{L})=0$. For $0<q<4$, a straightforward computation shows that the maximum bottleneck distance achieved is for $q=2$, where the cost is $2$. However, for the line $L^{\prime}$ through the points $(0,0)$ and $(7,11)$, the bottleneck distance is $d_{B}(M^{L},N^{L})=4$ and $\hat{m}^{L^{\prime}}=\frac{7}{11}$, so that the matching distance between $M$ and $N$ is at least $\frac{28}{11}$. This example shows that diagonal lines do not suffice.
@@
Let $M$ be the rectangle decomposable module which is the direct sum of two rectangle modules, one with underlying rectangle $0,7)\times[0,8)$, the other one with underlying rectangle $0,7)\times[4,11)$. Let $N$ be the rectangle decomposable module sum of two rectangle modules whose underlying rectangles are $0,7)\times[4,8)$ and $0,7)\times[0,11)$.
@@
Figure 6: Weighted bottleneck distance for the modules of Example 2.14 across the line parameters. On the horizontal axis we have the values of the angle $\theta$, giving the slope of the line. On the vertical axis we have the parameter $b$ giving the origin of the line. The red cross corresponds to the maximum value over all lines computed in the approximation, the black cross the maximum value for lines through pairs of critical values and the blue cross is the maximum value for diagonal lines. This particular example shows that it is not sufficient to consider only lines through pairs of critical values or diagonal lines. This is because the matching distance must be at least as much as the value computed at the red cross.
The following considerations confirms this: for diagonal lines $L$ with equation $y=x+q$ for $q\leq 1$ or $q\geq 4$, we have that $M^{L}=N^{L}$ so that $d_{B}(M^{L},N^{L})=0$. For $1<q<4$, straightforward computation shows that the maximum bottleneck distance achieved is for $q=2.5$, where the cost is $\frac{3}{2}$. However, for the line $L^{\prime}$ through the points $(\frac{7}{2},6)$ and $(7,11)$, the bottleneck distance is $d_{B}(M^{L},N^{L})=3$ and $\hat{m}^{L^{\prime}}=\frac{7}{10}$, so that the matching distance between $M$ and $N$ is at least $\frac{21}{10}$. Out of the lines through two points in the closure with respect to lowest upper bound of $C_{M}\cup C_{N}$, the only line that does not give $M^{L}=N^{L}$ is the line through $(0,0)$ and $(7,11)$, which has weighted bottleneck distance $\frac{21}{11}$. This example shows that it does not suffice to consider only diagonal lines and lines through two points in the closure with respect to the lowest upper bound of $C_{M}\cup C_{N}$.